Board? Games, Inc. makes board games. The following data pertains to the last six? months: Direct Labor Hours Manufacturing Overhead Month 1 ?45,000 ?$295,000 Month 2 67 comma 000 $ 317 comma 000 Month 3 ?57,000 ?$323,000 Month 4 ?52,000 ?$247,250 Month 5 ?34,000 ?$178,200 Month 6 25 comma 000 $ 165 comma 500 Based on this? data, what is the linear cost? equation? (Round intermediary calculations to the nearest? cent.) A. y? = $ 5.18X ?+ $ 34 comma 755 B. y? = $ 0.28X ?+ $ 151 comma 500 C. y? = $ 3.61X ?+ $ 75 comma 130 D. Cannot be determined from information given.?
Answer #1
Unit variable cost = (317000-165500)/(67000-25000)= 3.61
Fixed cost = 317000-(67000*3.61)= $75130
Option C $ 3.61X ?+ $ 75130 is correct